# DMesh++: An Efficient Differentiable Mesh for Complex Shapes

### v1

@@ Abstract
Recent probabilistic methods for 3D triangular meshes capture diverse shapes by differentiable mesh connectivity, but face high computational costs with increased shape details. We introduce a new differentiable mesh processing method in 2D and 3D that addresses this challenge and efficiently handles meshes with intricate structures. Additionally, we present an algorithm that adapts the mesh resolution to local geometry in 2D for efficient representation. We demonstrate the effectiveness of our approach on 2D point cloud and 3D multi-view reconstruction tasks. Visit our project page (https://sonsang.github.io/dmesh2-project) for source code and supplementary material.
Figure 1: DMesh++ for complex 2D and 3D shapes. (a) DMesh++ can reconstruct complex 2D drawings from sample points extracted from them. (b) In 2D, DMesh++ can produce efficient yet accurate meshes that adapt to local geometry. (c) DMesh++ can also reconstruct complex 3D shapes with colors from point clouds and multi-view images. For each result, the “imaginary” part is rendered in gray, while the “real” part, which defines the final mesh, is rendered in other colors (Sec. 3.1). In (b), the tone of the color represents the edge length.
* This work was done during internship at Adobe Research.
** The paper was last modified on Dec. 20th, 2024.
@@ 1 Introduction
Among various possible shape representations, a mesh is often favored for a wide range of downstream tasks due to its efficiency, versatility, and controllability. A mesh is defined by its vertices’ position and their connectivity in the form of edges and faces. This connectivity is discrete in nature, and also the number of possible connectivities grows exponentially with the number of points, which prevents meshes from being fully differentiable shape representations. To address this, recent data-driven efforts have attempted to predict mesh connectivity using Transformer-based models [36, 5, 6, 35, 11]. However, these methods face inherent challenges with robustness to outlier meshes, potential self-intersections, and high computational costs.
On another route, Son et al. [37] introduced a new form of differentiable mesh called DMesh, which is essentially a probabilistic approach. For a given set of points, they explicitly compute probabilities for possible face combinations to exist on the mesh. This approach guarantees mesh without self-intersections, and is free from outliers, as it is not data-driven. Therefore, this probabilistic approach opens up a new venue to adopt meshes in a machine learning pipeline, such as generative models [41, 45]. However, it suffers from excessive computational cost when the number of points increases (Fig. 6), which limits its applicability for representing complex shapes with detailed structures.
In this work, we introduce DMesh++, which overcomes the computational limitations of DMesh while retaining its core advantages. To that end, we present Minimum-Ball algorithm. This algorithm uses minimum circumscribing ball of a face to compute its probability. While the computational cost to evaluate face probability is $O(N)$ for DMesh, where $N$ is the number of points that define the mesh, our Minimum-Ball algorithm has $O(\log N)$ computational cost under practical scenarios (Secs. 3.2 and 6).
The direct application of DMesh++ is a reconstruction task. Using Minimum-Ball algorithm, we can reconstruct complex 2D and 3D meshes (Figs. 1, 7 and 11) efficiently within a reasonable time-frame. At the same time, as the reconstructed mesh usually has lots of redundant faces, we propose Reinforce-Ball algorithm to remove them, and produce an efficient mesh for 2D cases (Sec. 4.2). It is designed to optimize the probability of a point’s existence, and minimize the number of points while keeping the geometric accuracy (Tabs. 1, 1 and 8). Additionally, we propose several other novelties in reconstruction process (Secs. 4.3 and 4.4) for better reconstruction.
@@
  • We present DMesh++, an enhanced version of DMesh [37] that addresses its computational bottlenecks with Minimum-Ball algorithm.
  • To eliminate redundant faces and produce an efficient mesh that adapts to local geometry in 2D reconstruction, we introduce Reinforce-Ball algorithm.
  • We propose additional mesh operations to use in the reconstruction pipeline. We also present an improved differentiable renderer for 3D multi-view reconstruction.
  • Utilizing these components, we reconstruct high-quality meshes of complex shapes in both 2D and 3D, demonstrating our approach’s effectiveness in 2D point cloud reconstruction and 3D multi-view reconstruction tasks.
@@ 2 Related Work
While meshes offer an efficient and flexible representation of shapes, they are mainly constrained by their connectivity issues, which limit their applicability in machine learning. To address these challenges, shape inference in machine learning has evolved through three stages. We briefly introduce them below.
Using Alternative Differentiable Shape Representations. Rather than handling mesh directly, some prior work extract mesh from alternative differentiable shape representations. Neural implicit representations, like signed or unsigned distance fields [30, 43, 38, 39, 28, 40, 18, 20, 44], encode distance fields in neural networks, and use iso-surface extraction methods [21, 13, 10] to generate the mesh. Another method encodes distance directly into spatial points and applies differentiable iso-surface extraction [16, 33, 34, 40, 24, 19, 22]. While often more efficient, these methods typically cannot handle open surfaces; though [19] does, it cannot represent non-orientable geometries. Gaussian Splatting [14] also encodes visual data as spatial “splats” but lacks the geometric accuracy of implicit functions.
Inferring Meshes Differentiably. The main challenge in differentiable mesh handling is the exponential growth of possible vertex connections as vertex count increases. To simplify this challenge, most prior works assume (almost) fixed mesh connectivity [47, 4, 27, 17, 15, 29]. Recently, data-driven approaches have aimed to overcome these limitations by training generative models [36, 5, 6, 35, 11] that predict vertex connectivity from point clouds. Specifically, SpaceMesh [35] ensures combinatorial manifold mesh generation. However, these models still struggle with outliers and are susceptible to self-intersections.
Designing A Differentiable Form of Mesh. Son et al. [37] recently introduced DMesh, a differentiable mesh formulation using a probabilistic approach. DMesh augments each point with two continuous values, along with its position, and applies a “tessellation” function in Eq. 1 to deterministically generate a mesh from a point set. This method adapts to various geometric topologies and avoids self-intersections. With DMesh, optimizing or inferring only point-wise features is sufficient to generate the mesh, simplifying the application of losses on points or faces.
However, DMesh’s tessellation function is slow due to its reliance on Weighted Delaunay Triangulation (WDT), which has a practical time complexity of $O(N)$ for $N$ points using the CGAL package [12]. For $N=100K$ in 3D, the runtime can reach up to 800 milliseconds (Fig. 6), limiting DMesh’s applicability for complex shapes requiring finer detail. In this work, we eliminate WDT, proposing a more efficient differentiable mesh formulation.
@@ 3 Formulation
In this section, we first provide the high-level formulation for computing probability of a face to exist in the mesh. Then, we introduce Minimum-Ball (Sec. 3.2), which is one of our primary algorithms.
@@ 3.1 Preliminary
DMesh++ introduces an alternative tessellation function for faster processing. By removing the need for WDT, we eliminate the WDT weight and represent each point as a $(d+1)$-dimensional vector, $(x_{1},...,x_{d},\psi)$. In place of WDT, we implement a faster scheme called the Minimum-Ball condition (Definition 3.1) for defining the tessellation function. Letting $\mathbb{F}_{min}$ represent the set of faces that meet this condition, we define the tessellation function as
@@ 3.2 Minimum-Ball Algorithm
Figure 4: Reconstruction pipeline for 2D point cloud (up) and 3D multi-view images (down). Color represents per-point feature that is optimized: position, real ($\psi$), and probability ($\phi$).
The mesh generated by DMesh’s tessellation function in Eq. 1 is free from self-intersections, because $\mathbb{F}_{wdt}$ itself is free from it. However, the tessellation function is computationally expensive, as we need to compute WDT to define $\mathbb{F}_{wdt}$. In designing our Minimum-Ball, we desire to remove this necessity for acceleration, while inheriting the nice property about self-intersection. In the following, we show that Minimum-Ball satisfies these requirements.
@@
Note that we can ignore points in $F$, as they are located on the boundary of $B_{F}$. In Fig. 3, we render a 2D case, where $\overline{AB}$ does not satisfy this definition because of $F$. In contrast, $\overline{DF}$ satisfies this condition. Then, we can observe that $\mathbb{F}_{min}$ is a subset of faces in Delaunay Triangulation (DT) of $\mathbb{P}$, which we denote as $\mathbb{F}_{dt}$.
@@
Note that $\mathbb{F}_{dt}$ is also free from self-intersections as $\mathbb{F}_{wdt}$, and thus is $\mathbb{F}_{min}$. However, note that $\mathbb{F}_{min}$ does not necessarily tessellate the entire convex shape, as there could be faces in $\mathbb{F}_{dt}$ that are not in $\mathbb{F}_{min}$ (e.g. $\overline{AB}$ in Fig. 3).
Now, based on Definition 3.1, we can check if $F$ is in $\mathbb{F}_{min}$. Let us denote the center and radius of $B_{F}$ as $B^{c}_{F}\in\mathbb{R}^{d}$ and $B^{r}_{F}$. We can compute these values in a differentiable way (Appendix 7.2). Then, we can compute the maximum signed distance between $B_{F}$ and $\mathbb{P}$ as follows:
$d(B_{F},\mathbb{P})$ $=\max_{p\in\mathbb{P}-F}B^{r}_{F}-||p-B^{c}_{F}||$ (3)
$=B^{r}_{F}-\min_{p\in\mathbb{P}-F}||p-B^{c}_{F}||.$ (4)
@@
Then, it is straightforward to define $\Lambda_{min}$ with sigmoid function as
@@
4 Reconstruction
In this section, we discuss how to reconstruct DMesh++ in 2D and 3D from point clouds or multi-view images, as reconstruction is an immediate application of DMesh++. We introduce the reconstruction pipeline and novel algorithms to enhance mesh quality, including our second main algorithm, Reinforce-Ball. We also provide an overview of relevant loss functions.
4.1 Pipeline Overview
The goal of reconstruction is to optimize point-wise features, so that the resulting mesh aligns with the input observation – either a point cloud (2D) or multi-view images (3D) as shown in Figure 4. For implementation details about each step, please refer Appendix 8.
For 2D point cloud, we initialize points on a fixed triangular grid and set the “real” values based on their overlap with the point cloud (Step 1). We then fix the “real” values and optimize only the point positions (Step 2), using the Minimum-Ball algorithm to compute face existence probabilities (Sec. 3.2). To achieve an efficient yet geometrically accurate mesh, we use a novel Reinforce-Ball algorithm (Sec. 4.2) in Step 3.
For multi-view images, we initialize the 3D points on a tetrahedral grid and set the “real” values based on their projections onto the images (Step 1). As before, we fix the “real” values initially, optimizing only the point positions based on Minimum-Ball (Step 2), then proceed to optimize the “real” values again with fixed point positions (Step 3). Then, if desired, we remove non-manifoldness at this step. To achieve higher resolution, we apply a face subdivision scheme between epochs (Step 4). We propose algorithms for these mesh operations later (Sec. 4.3).
4.2 Reinforce-Ball Algorithm
Although the Minimum-Ball approach enables high-resolution mesh handling, it often results in overly dense meshes with redundant faces after reconstruction (Fig. 4). This occurs because Minimum-Ball lacks mesh complexity regularization. To address this in 2D cases, we propose Reinforce-Ball algorithm, which reduces unnecessary faces while preserving geometric detail.
Previously, DMesh [37] used “weight regularization” to simplify meshes by increasing regularization strength. However, this approach struggled with adaptive resolution due to local minima (Fig. 8). For example, when reconstructing the left line segment in inset, only the end points (black) are necessary, while the intermediate point (green) and colored edges are redundant. The bottom configuration is ideal, so if optimization starts from the top configuration, weight regularization reduces the redundant point’s weight, as shown in the middle configuration. However, this also lowers the probabilities of redundant edges, increasing reconstruction loss and thus restoring the redundant point’s weight, preventing convergence to the optimal configuration (Appendix 9.2). This issue arises from the combinatorial nature of the problem.
To solve this problem, our Reinforce-Ball algorithm defines per-point existence probability and optimize it using stochastic optimization technique [42]. The overview of this algorithm and its formal definition is given in Fig. 5 and Appendix 9.1, respectively.
$L^{i}_{rl}=L^{i}_{recon}+\epsilon_{card}\cdot L^{i}_{card},$ (8)
where $\epsilon_{card}$ is a small tunable hyperparameter to adjust the weight of the cardinality loss. We report ablation on it in Tabs. 1 and 8. If we write the final loss for a set of sampled points $\mathbf{P}$ as $L_{rl}(\mathbf{P})$, we aim at minimizing the expected loss:
We can estimate the gradient of this expected loss using log-derivative trick [25]. Then, we can update $\Phi(\mathbb{P})$ using this gradient, and remove unnecessary points while preserving the original geometry (Figs. 1, 4 and 8). Notably, this algorithm is particularly effective in 2D but less so in 3D, as explained in Appendix 9.3.
4.3 Mesh Operations
Here we propose two mesh operations that we use in the 3D reconstruction pipeline, which are easily applicable to 2D.
Non-Manifoldness Removal
Mesh defined by our tessellation function (Eq. 2) is not free from non-manifold edges or vertices in combinatorial sense. We remove non-manifold edges and vertices in Step 4. Non-manifold edges (inset, left) are edges that are adjacent to more than 2 faces, and non-manifold vertices (inset, right) are vertices whose adjacent faces do not form single fan. We detect such cases, and remove faces that are adjacent to such non-manifoldness until all of such cases are resolved, based on their contribution to the reconstruction loss (e.g. number of rendered pixels per face).
Mesh Subdivision
For mesh subdivision, we add points with $\psi=1$ at the middle of edges that are adjacent to currently existing faces. In the inset, points $E,D,F$ are the newly inserted points. They form 4 sub-faces with $A,B,C$, and they all satisfy Definition 3.1. Therefore, we can guarantee that these sub-faces will exist at the start of next epoch (Appendix 10).
4.4 Loss Functions
For 2D point cloud reconstruction, we use Chamfer Distance (CD) as the reconstruction loss. For 3D multi-view reconstruction task, we use rendering loss. Note that we need to formulate such loss functions to consider face probabilities ($\Lambda$Sec. 3.1) in Step 1, 2, and 4. We follow formulations of DMesh [37] – for CD, they propose expected CD, and for rendering, they interpret face probability as its opacity in their own differentiable renderer. However, their differentiable renderer lacks visibility gradients, and has to rely on another rendering process for them. To amend this issue, we improved their renderer by implementing (face) anti-aliasing in CUDA. As shown in the inset, we compute the intersecting area between each pixel and triangle to compute the face opacity at the pixel (Appendix 11). Finally, we use triangle quality regularization [37] to improve the mesh quality.

### v2

@@ Abstract
Recent probabilistic methods for 3D triangular meshes capture diverse shapes by differentiable mesh connectivity, but face high computational costs with increased shape details. We introduce a new differentiable mesh processing method that addresses this challenge and efficiently handles meshes with intricate structures. Our method reduces time complexity from $O(N)$ to $O(\log N)$ and requires significantly less memory than previous approaches. Building on this innovation, we present a reconstruction algorithm capable of generating complex 2D and 3D shapes from point clouds or multi-view images. Visit our project page for source code and supplementary material.
Figure 1: DMesh++ for complex 2D and 3D shapes. DMesh++ encodes all geometric and topological information into continuous point features. (a) By optimizing these point features, DMesh++ is able to reconstruct complex 2D drawings from sample points. (b) This approach is also applicable to 3D, where it reconstructs the complex geometric structure of DNA from a point cloud. (c) By incorporating additional color features, DMesh++ can reconstruct complex, colored 3D shapes from multi-view images. For each result, the “imaginary” part is rendered in gray, while the “real” part—which defines the final mesh—is rendered in other colors. (Sec. 3.1).
* This work was mainly done during internship at Adobe Research and continued as a collaborative effort with UMD.
** The paper was last modified on Jul. 6, 2025.
@@ 1 Introduction
Among various possible shape representations, a mesh is often favored for a wide range of downstream tasks due to its efficiency, versatility, and controllability. A mesh is defined by its vertices’ position and their connectivity in the form of edges and faces. This connectivity is discrete in nature, and also the number of possible connectivities grows exponentially with the number of points, which prevents meshes from being differentiable shape representations (Fig. 2). To address this, recent data-driven efforts have attempted to predict mesh connectivity using Transformer-based models [41, 5, 6, 40]. However, these methods face inherent challenges with robustness to outlier meshes, potential self-intersections, and high computational costs.
On another route, Son et al. [42] introduced a new form of differentiable mesh called DMesh, which is essentially a probabilistic approach. For a given set of points, they explicitly compute probabilities for possible face combinations to exist on the mesh based on the continuous point-wise features. This approach minimizes several mesh degeneracies, and is free from outliers, as it is not data-driven (Fig. 2). Therefore, this probabilistic approach opens up a new venue to adopt meshes in a machine learning pipeline, such as generative models [47, 51]. However, it suffers from excessive computational cost when the number of points increases (Fig. 7), which limits its applicability for representing complex shapes with detailed structures.
In this work, we introduce DMesh++, which overcomes the computational limitations of DMesh while retaining its core advantages. To that end, we present Minimum-Ball algorithm. While the computational cost to evaluate face probability is $O(N)$ for DMesh, where $N$ is the number of points that define the mesh, our Minimum-Ball algorithm has $O(\log N)$ computational cost (Secs. 3.2 and 7).
The direct application of DMesh++ is a reconstruction task. It effectively reconstructs complex 2D and 3D meshes of diverse topology from point clouds or multi-view images (Figs. 1, 9 and 10). During the optimization of continuous point-wise features, we observe dynamic topological changes in the mesh that recover the target shape (Fig. 3).
Figure 2: Conceptual comparison of traditional mesh and variants of DMesh [42]. Traditional meshes employ a non-differentiable, discrete data structure, $\mathbb{F}$, to store vertex indices that define connectivity, whereas DMesh++ encodes connectivity and additional information into continuous point-wise features, $\mathbb{P}$. Mesh generated from DMesh++ avoids several degeneracies—such as self-intersections and thin triangles—that can compromise its suitability for downstream applications.
@@
  • We present DMesh++, an enhanced version of DMesh [42], which overcomes its computational bottlenecks by employing the Minimum-Ball algorithm.
  • We propose a reconstruction algorithm that incorporates efficient loss formulations and additional mesh operations to effectively recover 2D and 3D shapes from point clouds or multi-view images.
  • We validate our approach on 500+ mesh models with diverse topology, which are collected from Thingi10K [52] and Objaverse [10] dataset.
@@ 2 Related Work
While meshes offer an efficient and flexible representation of shapes, they are mainly constrained by their connectivity issues, which limit their applicability in machine learning. To address these challenges, shape inference in machine learning has evolved through three stages.
Using Alternative Differentiable Shape Representations. Rather than handling mesh directly, some prior work extract mesh from alternative differentiable shape representations. Neural implicit representations, like (un)signed distance fields [34, 49, 44, 45, 32, 46, 22, 24, 50], encode distance fields in neural networks, and use iso-surface extraction algorithms [25, 16, 12] to generate the final mesh. Another method encodes distance directly into spatial points and applies differentiable iso-surface extraction [20, 38, 39, 46, 29, 23, 28]. While often more efficient, these methods typically cannot handle open surfaces; though [23] does, it cannot represent non-orientable geometries. Gaussian Splatting [18] also encodes visual data as spatial “splats” but lacks the geometric accuracy of implicit functions.
Figure 3: Mesh topology change during 3D point cloud reconstruction of a vase. We visualize the evolution of point-wise features by projecting them onto a 2-dimensional space using T-SNE [43] at each optimization step (ranging from 0 to 2K). As these features continuously evolve, the mesh undergoes discrete topological changes to progressively recover the target shape.
Inferring Meshes Differentiably. The main challenge in differentiable mesh handling is the exponential growth of possible vertex connections as vertex count increases. To simplify this challenge, most prior works assume fixed registration and permit only local connectivity changes [53, 4, 31, 21, 19, 33]. Recently, data-driven approaches have aimed to overcome these limitations by training generative models [41, 5, 6, 40] that predict vertex connectivity from point clouds. Specifically, SpaceMesh [40] ensures combinatorial manifold mesh generation. However, these models struggle with outliers and self-intersections.
Designing A Differentiable Form of Mesh. Son et al. [42] recently introduced DMesh, a differentiable mesh formulation using a probabilistic approach. DMesh augments each point with two continuous values, along with its position (Fig. 2), and applies a “tessellation” function in Eq. 1 to deterministically generate a mesh from a point set. This method adapts to various geometric topologies, including non-orientable open surfaces, and avoids self-intersections. With DMesh, optimizing or inferring only point-wise features is sufficient to generate the mesh.
However, DMesh’s tessellation function is slow due to its reliance on Weighted Delaunay Triangulation (WDT), which has a practical time complexity of $O(N)$ for $N$ points using the CGAL package [14]. For $N=100K$ in 3D, the runtime can reach up to 800 milliseconds (Fig. 7), limiting DMesh’s applicability for complex shapes requiring finer detail. Moreover, it is hardly possible to accelerate WDT using parallelization, because of its inherent race conditions. Therefore, in this work, we eliminate WDT, and propose a more efficient differentiable mesh formulation.
@@ 3 Formulation
In this section, we first provide the high-level formulation for computing probability of a face to exist in the mesh. Then, we introduce Minimum-Ball (Sec. 3.2), which is our primary algorithm.
@@ 3.1 Preliminary
DMesh++ introduces an alternative tessellation function for faster processing. By removing the need for WDT, we eliminate the WDT weight and represent each point as a $(d+1)$-dimensional vector, $(x_{1},...,x_{d},\psi)$ [^1] Points could carry additional features, such as color (Fig. 2).. In place of WDT, we implement a faster scheme called the Minimum-Ball condition (Definition 3.1) for defining the tessellation function. Letting $\mathbb{F}_{min}$ represent the set of faces that meet this condition, we define the tessellation function as
@@ 3.2 Minimum-Ball Algorithm
Figure 6: Reconstruction process for 3D multi-view colored images of a sculpture. In each stage, we optimize different per-point features: the position and the real ($\psi$), while the per-point color is refined at every stage. (Left) We display the meshes at each stage during the first epoch. (Right) We provide a rendering of the final mesh along with its view-point normal, which reveals the fine mesh details.
The mesh generated by DMesh’s tessellation function in Eq. 1 is free from self-intersections because $\mathbb{F}_{wdt}$ itself is free from them. Additionally, it minimizes the occurrence of thin triangles—undesirable in many downstream tasks such as physics simulations—thanks to the properties of the WDT. However, computing the tessellation function is computationally expensive, as it requires calculating the WDT to define $\mathbb{F}_{wdt}$. In designing our Minimum-Ball, we aimed to eliminate this computational bottleneck while preserving the favorable properties related to self-intersection avoidance and triangle quality. In the following, we demonstrate that Minimum-Ball satisfies these requirements.
@@
Note that we can ignore points in $F$, as they are located on the boundary of $B_{F}$. In Fig. 5, we render a 2D case, where $\overline{AB}$ does not satisfy this definition because of $F$. In contrast, $\overline{DF}$ satisfies this condition. Then, we can observe that $\mathbb{F}_{min}$ is a subset of faces in Delaunay Triangulation (DT) of $\mathbb{P}$ ($\mathbb{F}_{dt}$).
@@
Note that $\mathbb{F}_{dt}$ is also free from self-intersections as $\mathbb{F}_{wdt}$, and thus is $\mathbb{F}_{min}$. Furthermore, it inherently minimizes the number of thin triangles, as guaranteed by DT. However, note that $\mathbb{F}_{min}$ does not necessarily tessellate the entire convex shape, as there could be faces in $\mathbb{F}_{dt}$ that are not in $\mathbb{F}_{min}$ (e.g. $\overline{AB}$ in Fig. 5).
Now, based on Definition 3.1, we can check if $F$ is in $\mathbb{F}_{min}$. Let us denote the center and radius of $B_{F}$ as $B^{c}_{F}\in\mathbb{R}^{d}$ and $B^{r}_{F}$. We can compute these values in a differentiable way (Appendix 7.2). Then, we can compute the signed distance between $B_{F}$ and $\mathbb{P}$ as follows:
$d(B_{F},\mathbb{P})$ $=\min_{p\in\mathbb{P}-F}||p-B^{c}_{F}||-B^{r}_{F}.$ (3)
@@
Then, we define $\Lambda_{min}$ with sigmoid function as
@@
4 Reconstruction Process
The goal of reconstruction is to optimize point-wise features so that the resulting mesh aligns with the input observation. As shown in Fig. 3, the discrete mesh topology dynamically changes during the optimization of continuous features to better fit the given input.
Fig. 6 provides an overview of our reconstruction process for recovering a 3D colored mesh from multi-view images. In the first stage, we initialize per-point features. If a point cloud is available, we use it to obtain a better starting point (see the initial mesh in Fig. 3); otherwise, we initialize with a regular tetrahedral grid. Next, we optimize the point positions while keeping the real values fixed, followed by optimizing the real values while fixing the point positions. In these two stages, we minimize loss functions tailored to each input modality (e.g., Chamfer Distance loss for point clouds, rendering loss for images). To increase mesh complexity and capture finer details, we subdivide the mesh by inserting additional points. This process is iterated for a fixed number of epochs. Finally, to accelerate the overall process and enhance the final mesh quality, we introduce several innovations at each stage. Detailed explanations are provided in Appendix 8, due to lack of space.